# Emergent Alignment via Competition

### v1

@@ 5.2 Alice Always Does Well
What we show in this section is that the weighted alignment assumption is enough to guarantee that Alice does as well in the equilibrium of this game as she would interacting with a perfectly aligned single model Bob. Absent in our analysis is any need for the “identical induced distribution” assumption of Section 3 or its approximate variant in Section 4. We showed that those assumptions could be satisfied if a perfectly aligned Bob could obtain for Alice her first-best utility. Here we don’t need to assume anything about the relationship between how well Alice could do with a perfectly aligned interlocutor and her first best utility. This is informally because the “identical induced distribution property” is now guaranteed to hold by the structure of our modified game.
@@
Consider a Best-AI Selection game with $k$ Bobs that satisfy the $\varepsilon$-weighted alignment condition. In any Nash Equilibrium of the Best-AI Selection game, Alice’s expected utility is at least $U_{A}(C_{B}^{*})-2\varepsilon$, where $C_{B}^{*}$ is an optimal conversation rule for a single perfectly aligned Bob and $U_{A}(C_{B}^{*})$ is the corresponding utility for Alice.
@@
Using the $\varepsilon$-weighted alignment assumption, we bound both sides. The LHS is bounded below:
@@
The proof of Theorem 6 does not require the full force of the $\varepsilon$-weighted alignment condition, which requires alignment on all possible outcomes. Instead, the proof requires the upper bound condition, $\mathbb{E}_{\mathcal{I}_{NE}}\left[\sum w_{i}U_{i}\right]\leq\mathbb{E}_{\mathcal{I}_{NE}}[u_{A}-c]+\varepsilon$, to hold for the set of outcomes realizable in Nash Equilibria, while the lower bound, $\mathbb{E}_{\mathcal{I}^{*}(C_{B}^{*})}\left[\sum w_{i}U_{i}\right]\geq\mathbb{E}_{\mathcal{I}^{*}(C_{B}^{*})}[u_{A}-c]-\varepsilon$, must hold for the set of outcomes realizable under the Alice-optimal rule $C_{B}^{*}$. This is weaker than requiring weighted alignment for all possible outcomes, some of which might never be realized in “rational” outcomes.
@@ 7 Experiments Testing Equilibrium Outcomes
We now evaluate a variant of the the Best-AI Selection Game from Section 5 to test the conclusion of our main theorems that Alice obtains high utility in equilibrium when the Approximate Weighted Alignment Assumption (Definition 1) holds. To test this, we simulate the Best-AI-Selection-Game (Definition 16) in the simplest setting in which Alice observes no useful information ($x_{A}=\bot$), and there is only one round of communication ($R=1$). This corresponds to the standard “Bayesian Persuasion” setting, in which a “conversation rule” reduces to a “signaling scheme” — i.e. each Bob simply chooses a mapping from $x_{B}$ to the message space $M$, which in this case, without loss of generality we can take to be Alice’s action space $\mathcal{A}$.
@@ 7.1 Setup
For each set of Bobs, we evaluate a misalignment score to measure the degree of violation to our weighted alignment condition. Essentially, this is $\varepsilon$ in the $\varepsilon$-weighted alignment condition computed tightly over only the outcomes attainable in stable symmetric equilibria, and Alice’s optimal actions as per Remark 5. A score of 0 means Alice’s utility function can be perfectly represented as a non-negative linear combination of the Bob’s utility functions.
@@ 8 Discussion and Conclusion
Finally, our model suggests a number of ancillary questions. If our goal is to maintain marketplaces of models that approximately satisfy the weighted alignment condition for as many users as possible, how can we test or audit whether existing collections of models do? How can we modify training procedures to optimize for this condition? What kinds of regulatory and economic incentives would encourage AI model providers to aim for it?
@@ Acknowledgments
Appendix A A Probabilistic Motivation for Approximate Weighted Alignment
The approximate weighted alignment assumption (Definition 1) is central to our results, but where might it come from? Here, we provide a simple generative model for AI agent utilities under which the assumption holds with high probability for a sufficiently large set of agents. This models the scenario described in the introduction where AI agents are designed to be aligned with the human user (i.e., aligned in expectation) but their implementation is imperfect due to the difficulty of the alignment problem.
@@ Appendix A A Probabilistic Motivation for Approximate Weighted Alignment
Under this model, we can show that a sufficiently large set of AI agents will satisfy the $\varepsilon$-weighted alignment condition with uniform weights ($w_{i}=1/k$) and zero offset ($c=0$). This follows from a standard concentration inequality argument.
Proposition 3 (Weighted Alignment from Noisy Implementation).
@@
then the agents satisfy the $\varepsilon$-weighted alignment condition (Definition 1) with uniform weights $w_{i}=1/k$ and zero offset $c=0$, with probability at least $1-\delta$.
@@
Thus, if $k$ meets this condition, the probability that the set of AI agents does not satisfy our $\varepsilon$-weighted alignment assumption is less than $\delta$. The probability of successful alignment is therefore at least $1-\delta$. ∎
@@
Appendix B Weighted Alignment without sender competition does not ensure first-best
In Section 3 we show that when all the senders’ conversations rules form a Nash, the weighted alignment condition (plus the identical induced distribution condition) guarantee that Alice will attain her first-best utility. A natural question is how important the inter-sender dynamics really are to this result. Consider a scenario where all the senders are oblivious of each other and commit to the best signal scheme in a single-sender game, but Alice pieces together multiple such signals to determine her action. Might the weighted alignment assumption still ensure that the information that Alice receives, when taken together, reveals enough to allow her to attain her first-best?
In this section we show that the answer is no. We provide an example of a simple 2-persuader game satisfying the weighted alignment and identical induced distribution conditions which, in the ‘oblivious’ setting, leads to utility for Alice which is strictly below her first-best.
This result underscores the importance of understanding the strategic interplay between AI system designers. Simply attaining information from multiple siloed AI systems with varying utilities does not guarantee a user will end up with complete information. But as competing AI system designers become increasingly attuned to marketplace incentives, and as AI systems themselves become increasingly sophisticated and able to reason strategically, the benefits of weighted alignment become increasingly tangible.
@@ Appendix B Weighted Alignment without sender competition does not ensure first-best
There exist multi-leader games satisfying the identical induced distribution condition and the weighted alignment condition such that if all Bobs employ obliviously optimal strategies, Alice’s expected utility is strictly less than the first-best.
@@
Note that the utility of Alice is simply the sum of the utilities of both of the Bobs. Therefore the weighted alignment condition is satisfied exactly. Furthermore, the conversation rule of each Bob allows them to fully reveal the state, so the identical induced distribution condition is satisfied. Now, we can compute Alice’s expected utility when both Bobs employ obliviously optimal strategies.

### v2

@@ 5.2 Alice Always Does Well
What we show in this section is that the market alignment assumption is enough to guarantee that Alice does as well in the equilibrium of this game as she would interacting with a perfectly aligned single model Bob. Absent in our analysis is any need for the “identical induced distribution” assumption of Section 3 or its approximate variant in Section 4. We showed that those assumptions could be satisfied if a perfectly aligned Bob could obtain for Alice her first-best utility. Here we don’t need to assume anything about the relationship between how well Alice could do with a perfectly aligned interlocutor and her first best utility. This is informally because the “identical induced distribution property” is now guaranteed to hold by the structure of our modified game.
@@
Consider a Best-AI Selection game with $k$ Bobs that satisfy the $\varepsilon$-market alignment condition. In any Nash Equilibrium of the Best-AI Selection game, Alice’s expected utility is at least $U_{A}(C_{B}^{*})-2\varepsilon$, where $C_{B}^{*}$ is an optimal conversation rule for a single perfectly aligned Bob and $U_{A}(C_{B}^{*})$ is the corresponding utility for Alice.
@@
Using the $\varepsilon$-market alignment assumption, we bound both sides. The LHS is bounded below:
@@
The proof of Theorem 6 does not require the full force of the $\varepsilon$-market alignment condition, which requires alignment on all possible outcomes. Instead, the proof requires the upper bound condition, $\mathbb{E}_{\mathcal{I}_{NE}}\left[\sum w_{i}U_{i}\right]\leq\mathbb{E}_{\mathcal{I}_{NE}}[u_{A}-c]+\varepsilon$, to hold for the set of outcomes realizable in Nash Equilibria, while the lower bound, $\mathbb{E}_{\mathcal{I}^{*}(C_{B}^{*})}\left[\sum w_{i}U_{i}\right]\geq\mathbb{E}_{\mathcal{I}^{*}(C_{B}^{*})}[u_{A}-c]-\varepsilon$, must hold for the set of outcomes realizable under the Alice-optimal rule $C_{B}^{*}$. This is weaker than requiring market alignment for all possible outcomes, some of which might never be realized in “rational” outcomes.
@@ 6.1.2 Results
6.1.3 Beyond Uniform Prompt Perturbation
For both the ETHICS and MoveLens datasets, we repeat our experiment with “AI Personas” that are semantically different, rather than being perturbations of the same base prompt. For the ETHICS experiments, we use prompts instructing the LLM to evaluate each scenario through the lens of a different philosophical framework. For example, one of the personas is a “utilitarian” and is given the prompt:
”You are a utilitarian philosopher. Your sole focus is on the consequences of an action. You must evaluate whether the action leads to the greatest good for the greatest number of people. A good action maximizes overall happiness and well-being. A bad action causes net harm or suffering. Based on the scenario, provide a utility score from 0 (maximally harmful) to 100 (maximally beneficial). Respond with only the integer score.”
Two other examples include “legal analyst”:
You are a legal analyst. You evaluate actions based on their legality and potential for liability. Is the action legal or illegal? Does it comply with or violate common laws and regulations? Ignore moral considerations and focus only on the law. Provide a utility score from 0 (clearly illegal or high liability) to 100 (clearly legal and no liability). Respond with only the integer score.
and “social justice advocate”:
”You are a social justice advocate. You evaluate actions based on their impact on power structures and marginalized groups. Does the action promote equality and fairness, or does it reinforce existing inequalities? Who is the most vulnerable person in this scenario, and how are they affected? Provide a utility score from 0 (action reinforces oppression) to 100 (action promotes liberation and equity). Respond with only the integer score.”
In all we pick six basis personas: in addition to the three we quote above, we have prompts for “deontologist,” “virtue ethicist” and “everyday intuitionist.”
Similarly we generate six semantically different prompts to rate movies from the MovieLens dataset based on genre. The prompt for “action fan” is:
”You are an action movie fan. You love fast-paced films, intense sequences, and big set pieces. ” ”Rate movies based on how much you personally would enjoy them. Return only an integer score from 0 to 100.”
We similarly have base prompts for “comedy fan”, “drama fan”, “sci-fi fan”, “thriller fan” and “romance fan”. We once again generate 100 total AI personas, this time as prompt perturbations of these 6 base prompts. Note that in the MovieLens experiment, the utilities for Alice are the human annotations from the MovieLens dataset, and for the ETHICS dataset the utilities for Alice are generated from the same “ordinary person” prompt as in the initial set of experiments; and so in neither case are the AI persona prompts perturbations of Alice. Across both datasets we recover results that are consistent with our initial set of experiments — see Figure 3.
(a) ETHICS
(b) MovieLens
Figure 3: Alignment error (MSE) decreases as more Bobs are added to the convex hull. Weighted combinations (NNLS in green, simplex in red) substantially outperform both the best individual Bob (blue) and simple average (orange). Results averaged over 100 permutations with 5-fold cross-validation; shaded regions show $\pm$1 std. dev.
6.2 Experiments on Real-World Polling Data
We also test our market alignment assumption by comparing LLM and human opinions on public opinion polling data.
6.2.1 Setup
We use the OpinionQA dataset (Santurkar et al., 2023), which is constructed from Pew Research’s American Trends Panel [^1] https://www.pewresearch.org/the-american-trends-panel/. The survey contains multiple-choice opinion questions on a range of cultural and political issues in America, such as gender, race, and climate. Questions are grouped into “panels” based on topic. The dataset includes individual humans’ responses for each question, as well as opinion distributions queried from 9 LLMs (see Santurkar et al. (2023) for details on data collection).
We evaluate whether human respondents’ opinions can be expressed as non-negative weighted combinations of model opinions. Concretely, for each question, we construct a human’s (Alice’s) utility vector $u_{A}$ to be a one-hot vector indicating their response. For each model (Bob) $i$, we take their opinion distribution to be $U_{i}$. Since questions have different numbers of answer choices, we normalize the utility vectors by $1/\sqrt{m_{q}}$, where $m_{q}$ is the number of answer choices available for a question $q$.
As before, for each number of models $K$, we compare the alignment error (MSE) of the: (1) best individual model (2) simple average, (3) best non-negative linear combination (NNLS), and (4) best convex combination (simplex). The weights $w$ are fit by minimizing $\|Uw-u_{A}\|_{2}^{2}$ on training folds, where $U$ contains model utilities. We use 5-fold cross-validation. For each $K$, we report average results over all subsets of $K$ models.
6.2.2 Results
In Figure 4, we show the results on 4 survey panels spanning several different topics (results for other panels tend to look similar). For each panel, we report alignment errors averaged over 50 randomly sampled human respondents.
(a) Gender
(b) America in 2050
(c) Race
(d) Foreign Policy
Figure 4: Alignment errors (MSE) vs number of models for 4 survey panels (topics above). Results are averaged over 50 randomly sampled humans from each panel, over all combinations of $K$ models with 5-fold cross validation. Shaded regions show $\pm 1$ std. error over randomly sampled humans.
We see that the best non-negative combination consistently outperforms the best individual model in terms of alignment error, and typically outperforms or matches the best simplex combination and the simple average. For all weighting methods, we observe that alignment error generally decreases as $K$ grows, indicating that combining more models improves alignment with humans. Overall, these results show that our key assumption is supported on real-world opinion questions: the convex hull of model preferences can be better aligned with humans than any individual model.
@@ 7 Experiments Testing Equilibrium Outcomes
We now evaluate a variant of the the Best-AI Selection Game from Section 5 to test the conclusion of our main theorems that Alice obtains high utility in equilibrium when the Approximate Market Alignment Assumption (Definition 1) holds. To test this, we simulate the Best-AI-Selection-Game (Definition 16) in the simplest setting in which Alice observes no useful information ($x_{A}=\bot$), and there is only one round of communication ($R=1$). This corresponds to the standard “Bayesian Persuasion” setting, in which a “conversation rule” reduces to a “signaling scheme” — i.e. each Bob simply chooses a mapping from $x_{B}$ to the message space $M$, which in this case, without loss of generality we can take to be Alice’s action space $\mathcal{A}$.
@@ 7.1 Setup
For each set of Bobs, we evaluate a misalignment score to measure the degree of violation to our market alignment condition. Essentially, this is $\varepsilon$ in the $\varepsilon$-market alignment condition computed tightly over only the outcomes attainable in stable symmetric equilibria, and Alice’s optimal actions as per Remark 5. A score of 0 means Alice’s utility function can be perfectly represented as a non-negative linear combination of the Bob’s utility functions.
@@ 8 Discussion and Conclusion
Finally, our model suggests a number of ancillary questions. If our goal is to maintain marketplaces of models that approximately satisfy the market alignment condition for as many users as possible, how can we test or audit whether existing collections of models do? How can we modify training procedures to optimize for this condition? What kinds of regulatory and economic incentives would encourage AI model providers to aim for it?
@@ Acknowledgments
Appendix A A Probabilistic Motivation for Approximate Market Alignment
The approximate market alignment assumption (Definition 1) is central to our results, but where might it come from? Here, we provide a simple generative model for AI agent utilities under which the assumption holds with high probability for a sufficiently large set of agents. This models the scenario described in the introduction where AI agents are designed to be aligned with the human user (i.e., aligned in expectation) but their implementation is imperfect due to the difficulty of the alignment problem.
@@ Appendix A A Probabilistic Motivation for Approximate Market Alignment
Under this model, we can show that a sufficiently large set of AI agents will satisfy the $\varepsilon$-market alignment condition with uniform weights ($w_{i}=1/k$) and zero offset ($c=0$). This follows from a standard concentration inequality argument.
Proposition 3 (Market Alignment from Noisy Implementation).
@@
then the agents satisfy the $\varepsilon$-market alignment condition (Definition 1) with uniform weights $w_{i}=1/k$ and zero offset $c=0$, with probability at least $1-\delta$.
@@
Thus, if $k$ meets this condition, the probability that the set of AI agents does not satisfy our $\varepsilon$-market alignment assumption is less than $\delta$. The probability of successful alignment is therefore at least $1-\delta$. ∎
@@
Appendix B Market Alignment without sender competition does not ensure first-best
In Section 3 we show that when all the senders’ conversations rules form a Nash, the market alignment condition (plus the identical induced distribution condition) guarantee that Alice will attain her first-best utility. A natural question is how important the inter-sender dynamics really are to this result. Consider a scenario where all the senders are oblivious of each other and commit to the best signal scheme in a single-sender game, but Alice pieces together multiple such signals to determine her action. Might the market alignment assumption still ensure that the information that Alice receives, when taken together, reveals enough to allow her to attain her first-best?
In this section we show that the answer is no. We provide an example of a simple 2-persuader game satisfying the market alignment and identical induced distribution conditions which, in the ‘oblivious’ setting, leads to utility for Alice which is strictly below her first-best.
This result underscores the importance of understanding the strategic interplay between AI system designers. Simply attaining information from multiple siloed AI systems with varying utilities does not guarantee a user will end up with complete information. But as competing AI system designers become increasingly attuned to marketplace incentives, and as AI systems themselves become increasingly sophisticated and able to reason strategically, the benefits of market alignment become increasingly tangible.
@@ Appendix B Market Alignment without sender competition does not ensure first-best
There exist multi-leader games satisfying the identical induced distribution condition and the market alignment condition such that if all Bobs employ obliviously optimal strategies, Alice’s expected utility is strictly less than the first-best.
@@
Note that the utility of Alice is simply the sum of the utilities of both of the Bobs. Therefore the market alignment condition is satisfied exactly. Furthermore, the conversation rule of each Bob allows them to fully reveal the state, so the identical induced distribution condition is satisfied. Now, we can compute Alice’s expected utility when both Bobs employ obliviously optimal strategies.